Write the equation of the line passing through (-2,-6) and (4,-6)
Found 2 solutions by scott8148, checkley75:
Answer by scott8148(6628)

the y-coordinate is not changing, so the equation is y=-6

Answer by checkley75(3666)

FIRST YOU NEED TO FIND THE SLOPE(m)
m=(Y2-Y1)/(X2-X1)
m=(-6+6)/(4+2)=0/6=0 OR THIS IS A HORIZONTAL LINE THROUGH [Y=-6] WHICH IS ALSO THE LINE EQUATION